OpenStudy (anonymous):

Solve for x: x2 + 24x + 90 = 0

OpenStudy (anonymous):

Use the quadratic formula.

OpenStudy (anonymous):

Would you know how to do that?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

If you do solve it, let me know what you get.

OpenStudy (ashleyisakitty):

x=-3(4+√6)

OpenStudy (anonymous):

i got x= -12+3√6, -12-3√6

OpenStudy (anonymous):

Is that right?

OpenStudy (ashleyisakitty):

You were damn close!

OpenStudy (anonymous):

I got a little different answer. x=3(-4-√6) and x=3(√6 -4)

OpenStudy (anonymous):

But when you multiply them out you get what i got